OpenStudy (sydney3545):

Solve: -4(2x - 1) = -6(x + 2) - 2 -5 5 9 -9

OpenStudy (anonymous):

Solving for x?

OpenStudy (sydney3545):

yes

OpenStudy (anonymous):

Okay. So, first we want to distribute which will get you \[-8x+4=(-6x-12)-2\]

OpenStudy (anonymous):

No what to do now?

OpenStudy (anonymous):

Know* sorry

OpenStudy (sydney3545):

No :/ thats why im stuck is at this point

OpenStudy (anonymous):

add 6x to your left side which gives you -2x +4 =-12 -2

OpenStudy (anonymous):

does that help at all?

OpenStudy (anonymous):

Oh. Sorry

OpenStudy (anonymous):

-2x=-18 which gives you positive 9

OpenStudy (xxdragoonxx):

-4(2x - 1) = -6(x + 2) - 2 = -4*2x-4*-1=6*x-6*2-2 = -8x+4=-6x-12-2 = -8x+4=-6x-14 = -8x+4+6x=-14 = -8x+6x+4=-14 = -2x+4=-14 = -2x=-14-4 = -2x=-18 x = -18/-2 x =9